# A new Angle on Benchmarking Noncovalent Interactions

### v2

@@ 3.8 Assessment of the results when extrapolating to larger molecules
How would this translate to other, similar systems and their intermolecular interactions, such as the coronene dimer from the L7 set of molecules and the buckycatcher? There are several aspects to this: Clearly, the interaction energy of the buckycatcher of approximately 175 kJ/mol at the CCSD(T) level of theory is much larger than the interaction energy of the hexacene dimer with our best estimate at approximately 61 kJ/mol. The latter, with six rings, is more comparable to the coronene dimer with seven rings, and an interaction energy of 76 kJ/mol (FN-DMC) to 86 kJ/mol (CCSD(T)). However, the small discrepancy between CCSD(T) and CCSDT(Q) is rather telling, as it is only 1.5 kJ/mol for the hexacene dimer. Even extrapolating the same slope to the buckycatcher, the deviation would only sum up to approximately 4 kJ/mol difference, indicating that the deviations between CCSDT(Q) and CCSD(T) for intermolecular interactions are an order of magnitude smaller than the 44 kJ/mol difference between FN-DMC and CCSD(T) of Table 1.
@@
If we do not compare interaction energies, but rather MP2-MP3 and MP2-CCSD differences of 128 (MP2-MP3) and 98 kJ/mol (MP2-CCSD) for the coronene dimer, and a (rather large) MP2-CCSD difference of 261.1 kJ/mol for the buckycatcher, we will arrive at different numbers by extrapolating to these values, considering the slopes obtained above for the acenes. This MP2-CCSD difference of the slope is 15.3 0.1 kJ/mol per acene ring at the basis set limit, taking cp-corrected and uncorrected values into account. Interestingly, the MP2-CCSD difference is highly basis set dependent, as for the cc-pVDZ basis set, the difference in the slope is only 10.8 0.5 kJ/mol per acene ring. The deviations in the slopes between CCSD(T) and CCSDT(Q) are, in comparison, only 0.36 0.02 kJ/mol, which is a factor of 40 smaller than the MP2-CCSD difference.
@@
Extrapolating this difference again to the buckycatcher, this would amount to a deviation of 5.6 kJ/mol (and 2.3 kJ/mol for the coronene dimer). Unfortunately, we have CCSDT(Q) values using only a double-zeta basis set. However, we can compute other approximate methods using a larger basis set, compare, and then scale the CCSDT(Q)-CCSD(T) differences in the slopes by these values. For the acenes, the MP2-CCSD(T) slopes differ by 21% when going from a double-zeta basis set to an extrapolated cc-pV{T,Q}Z value . In contrast, the CCSD-CCSD(T) slopes differ by 57% for the same basis sets (DZ and cc-pV{T,Q}Z). As we can see from Table 2, the CCSDT-2 slopes are rather close to CCSDT(Q), and the CCSDT-2-CCSD(T) slopes behave slightly more like CCSD-CCSD(T) for a small basis set of triple-zeta quality, see Supporting Information. Because of this, we assume that at the double-zeta basis set for the CCSD(T)-CCSDT(Q) slope the difference is underestimated by a maximum of 40%. Hence, we would obtain a deviation from CCSD(T) of 7.9 kJ/mol for the buckycatcher and 3.2 kJ/mol for the coronene dimer.
@@
The large deviations in the slopes of CCSD and MP2 in comparison to that of CCSD(T) and CCSDT(Q) also imply that a different angle when looking at intermolecular interactions is needed, and approximate methods may unfortunately not provide a full picture.
For the L7 systems and the buckycatcher dimer in question, the ethylene stacks are not as much of interest, as their post-CCSD(T) effects can be considered as negligible here, the CCSD(T) and CCSDT(Q) slopes are extremely similar, arriving at the CCSDT(Q) line y = (-8.7 0.2 kJ/mol)x + 6.5 0.2 kJ/mol per ethylene double bond. In comparison, the final best CCSDT(Q) lines obtained for the acenes are = (-10.5 0.2 kJ/mol) + 5.2 0.3 kJ/mol for the full set of the six acenes. The interaction energies of benzene () and naphthalene () sandwich structures are -6.5 0.2 kJ/mol and -15.5 0.2 kJ/mol, respectively: the resulting slope would be slightly different from the one discussed above because the two data points include the Hartree-Fock association energy, which has its own slope. The interaction energies of ethylene and trans-butadiene, within the geometric constraints applied, are -1.7 kJ/mol and -7.3 kJ/mol. These small molecules or the line of the series can serve as reference for other methods when testing the increase in interaction energy for such species.
@@ 4 Conclusions
In case we move to ever larger molecules, we propose a new angle of looking at intermolecular interactions of molecular dimers. Although local methods already provide a good, alternative avenue for calculating these, we are unfortunately still away from drawing definite conclusions about the accuracy of electronic structure methods, such as the gold CCSD(T) standard.
Here, we use different dimer series exhibiting very linear increases of the correlation energy, such as acene and ethylene stacks. We were able to show that none of the following factors disrupt this linear behavior: the choice of basis set; the local correlation cutoffs; the electronic structure method; fixed vs. relaxed geometry. This is obvious from all figures in this paper.
@@
Concerning the basis sets, cc-pVTZ already is sufficient to obtain acceptable slopes, with a deviation in the slopes of 1.5 kJ/mol per ring for the acene series. For local correlation methods, only the tightest cutoffs (vvTight or TCutPNO=7) achieve this accuracy, whereas other cutoffs can deviate quite significantly with more than 4 kJ/mol per ring.
Finally, we are able to estimate the interaction of large acenes and ethylene stacks which are of similar interaction energy size or MP2-CCSD difference than the L7 set of molecules by comparing CCSD(T) to CCSDT(Q) slopes: Here, the deviations are much smaller than anticipated, with the CCSDT(Q)-CCSD(T) difference in the slope of 0.4 kJ/mol per ring. For comparison, this value is smaller than the deviation of canonical CCSD(T) from local CCSD(T) with the tightest cutoffs or medium-size basis sets of (aug-)TZ quality from the basis set limit, showing that these two effects can still pose major challenges.
By converging all these effects on the slopes to 0.2 kJ/mol per ring or double bond, we can estimate deviations of interaction energies from CCSD(T) at the basis set limit for large acene and ethylene stacks, and ultimately very large molecules including those which are having huge polarizabilities.
In case of the acene dimer series, CCSD(T) exhibits an increased interaction energy of 3.4% in comparison to CCSDT(Q), which is sizeable, but nowhere near the level that would explain the discrepancies to FN-DMC. For the ethylene stacks, on the other hand, there is no notable difference between CCSD(T) and CCSDT(Q). For the ethylene stacks, on the other hand, there is no notable difference between CCSD(T) and CCSDT(Q). Thus, with some limitations, we can reconfirm the ”gold standard” of basis set limit CCSD(T) for intermolecular interactions.
@@ Acknowledgments
This research was supported by an internal grant to JMLM from the Uriel Arnon Memorial Fund for Artificial Intelligence in Materials Research, as well as by a very generous allocation of CPU time on the Weizmann Faculty of Chemistry HPC facility chemfarm. The authors thank its administrators, Dr. Mark Vilensy and Andrei Vasilev, for their kind assistance. ADB was supported by the Weizmann Institute of Science as a Weston Visiting Scholar.

### v3

@@ 3.8 Assessment of the results when extrapolating to larger molecules
How would this translate to other, similar systems and their intermolecular interactions, such as the coronene dimer from the L7 set of molecules and the buckycatcher? There are several aspects to this: Clearly, the interaction energy of the buckycatcher of approximately 175 kJ/mol at the CCSD(T) level of theory is much larger than the interaction energy of the hexacene ‘sandwich’ dimer with our best estimate at approximately 61 kJ/mol. The latter, with six rings, is more comparable to the coronene dimer with seven rings, with an interaction energy of 76 kJ/mol (FN-DMC) to 86 kJ/mol (CCSD(T)). However, the small discrepancy between CCSD(T) and CCSDT(Q) is rather telling, as it is only 1.5 kJ/mol for the hexacene dimer. Even extrapolating the same slope to the buckycatcher, the deviation would only sum up to approximately 4 kJ/mol difference, indicating that the deviations between CCSDT(Q) and CCSD(T) for intermolecular interactions are an order of magnitude smaller than the 44 kJ/mol difference between FN-DMC and CCSD(T) of Table 1.
@@
If we do not compare interaction energies, but rather MP2MP3 and MP2CCSD differences of 128 (MP2MP3) and 98 kJ/mol (MP2CCSD) for the coronene dimer, and a (rather large) MP2CCSD difference of 261 kJ/mol for the buckycatcher, we will arrive at different numbers by extrapolating to these values, considering the slopes obtained above for the acenes. This MP2CCSD difference of the slope is 15.3 0.1 kJ/mol per acene ring at the basis set limit, taking CP-corrected and uncorrected values into account. Interestingly, the MP2CCSD difference is highly basis set dependent, as for the DZ basis set, the difference in the slope is only 10.8 0.5 kJ/mol per acene ring. The deviations in the slopes between CCSD(T) and CCSDT(Q) are, in comparison, only 0.36 0.02 kJ/mol, which is a factor of 40 smaller than the MP2-CCSD difference.
@@
Extrapolating this difference again to the buckycatcher, this would amount to a deviation of 5.6 kJ/mol (and 2.3 kJ/mol for the coronene dimer). Unfortunately, we have CCSDT(Q) values using only a double-zeta basis set. However, we can compute energies at more approximate levels in a larger basis set, compare, and then scale the CCSDT(Q)-CCSD(T) differences in the slopes by these values. For the acenes, the MP2-CCSD(T) slopes differ by 20% when going from a double-zeta basis set to an extrapolated cc-pV{T,Q}Z value. In contrast, the CCSD-CCSD(T) slopes differ by 40% for the same basis sets (DZ and cc-pV{T,Q}Z), and the ethylene stacks yield similar results. As we can see from Table 2, the CCSDT-2 slopes are rather close to CCSDT(Q), and the CCSDT-2-CCSD(T) slopes behave slightly more like CCSD-CCSD(T) for a small basis set of triple-zeta quality, see Supporting Information. Because of this, we assume that, using the double-zeta basis set for the CCSD(T)-CCSDT(Q) slope, the difference is underestimated by a maximum of 40%. Hence, we would obtain a deviation from CCSD(T) of 7.9 kJ/mol for the buckycatcher and 3.2 kJ/mol for the coronene dimer.
@@
The large deviations in the slopes of CCSD and MP2 in comparison to that of CCSD(T) and CCSDT(Q) also imply that a different perspective when looking at intermolecular interactions is needed, and approximate methods may unfortunately not provide a full picture.
For the L7 systems and the buckycatcher dimer in question, the ethylene stacks are not of as much interest, as their post-CCSD(T) effects can be considered as negligible here. The CCSD(T) and CCSDT(Q) slopes are extremely similar, arriving at the CCSDT(Q) line including HF (not just the correlation energy): = (-3.3 0.1 kJ/mol) + 3.7 0.2 kJ/mol per ethylene double bond.
In comparison, the final best CCSDT(Q) lines obtained for the acenes (including HF) are = (-10.5 0.2 kJ/mol) + 5.2 0.3 kJ/mol for the full set of the six acenes. The interaction energies of benzene () and naphthalene () ‘sandwich’ structures are -6.5 0.2 kJ/mol and -15.5 0.2 kJ/mol, respectively: the resulting slope would be slightly different from the one discussed above because the two data points include the Hartree-Fock association energy, which has its own slope. The interaction energies of ethylene and trans-butadiene, within the geometric constraints applied, are -0.36 0.03 kJ/mol and -2.92 0.09 kJ/mol. These small molecules or the line of the series can serve as reference for other methods when testing the increase in interaction energy for such species.
@@ 4 Conclusions
In case we move to ever larger molecules, we propose a new perspective for analyzing intermolecular interactions of molecular dimers. Although local methods already provide a good, alternative avenue for calculating these interactions, we have unfortunately not quite reached the point where we can draw definite conclusions about the accuracy of electronic structure methods such as the CCSD(T) ‘gold standard’.
Here, we use different dimer series exhibiting very linear increases in correlation energy, such as acene and ethylene stacks. We were able to show that none of the following factors disrupt this linear behavior: the choice of basis set; the local correlation cutoffs; the electronic structure method; and fixed vs. relaxed geometry. This is evident from all figures in this paper.
@@
Regarding basis sets, cc-pVTZ is already sufficient to obtain acceptable slopes, with a deviation in the slopes of 1.5 kJ/mol per ring for the acene series. For local correlation methods, only the tightest cutoffs (vvTight or TCutPNO=7) achieve this accuracy, whereas other cutoffs can deviate quite significantly, by more than 4 kJ/mol per ring.
Finally, we are able to estimate the interaction of large acenes and ethylene stacks which are of similar interaction energy size or MP2-CCSD difference than the L7 set of molecules by comparing CCSD(T) to CCSDT(Q) slopes: Here, the deviations are much smaller than anticipated, with the CCSDT(Q)CCSD(T) difference in the slope of 0.4 kJ/mol per ring. For comparison, this value is smaller than the deviation of canonical CCSD(T) from local CCSD(T) with the tightest cutoffs or medium-size basis sets of (aug-)TZ quality from the basis set limit, showing that these two effects can still pose major challenges.
By converging all these effects on the slopes to 0.2 kJ/mol per ring or double bond, we can estimate deviations of interaction energies from CCSD(T) at the basis set limit for large acene and ethylene stacks, and ultimately for very large molecules including those which are having huge polarizabilities.
In the case of the acene dimer series, CCSD(T) exhibits an interaction energy excess of 3.4% compared to CCSDT(Q) — which is sizable, but nowhere near the level that would explain the discrepancies to FN-DMC. For the ethylene stacks, on the other hand, there is no noteworthy difference between CCSD(T) and CCSDT(Q). Thus, with some reservations, we can reaffirm the ‘gold standard’ of basis set limit CCSD(T) for intermolecular interactions.
@@ Acknowledgments
This research was supported by an internal grant to JMLM from the Uriel Arnon Memorial Fund for Artificial Intelligence in Materials Research, as well as by a very generous allocation of CPU time on the Weizmann Faculty of Chemistry HPC facility chemfarm. The authors thank its administrators, Dr. Mark Vilensy and Andrei Vasilev, for their kind assistance, and Dr. Margarita Shepelenko for critically reading the draft. ADB was supported by the Weizmann Institute of Science as a Weston Visiting Scholar.